OpenStudy (anonymous):

If A and B are independent events, P(A) = 0.25, and P(B) = 0.45, find the probabilities below. (Enter your answers to four decimal places.) (a) P(A ∩ B) (b) P(A ∪ B) (c) P(A | B) (d) P(Ac ∪ Bc)

Parth (parthkohli):

to get P(A union B), you multiply the probabilities.

OpenStudy (anonymous):

I got a and c, need help with b and d.

OpenStudy (anonymous):

b) P(A or B) = P(A) + P(B) - P(A and B) d) P(~A or ~B) = P(~ (A and B)) = 1 - P(A and B)

OpenStudy (anonymous):

Thanks! Still stuck on d.. how do I get P(~ (A and B))? @sourwing

OpenStudy (anonymous):

P(~ (A and B)) = 1 - P(A and B)

OpenStudy (anonymous):

Oh ok I see now, thanks!